OpenStudy (anonymous):

There is another question @Abhisar As the chain moves, it turns the small gear. If you use the distance you calculated in the previous question, through what angle does the small gear turn to the nearest degree? idk if you can help..........if you cant its fine tho

OpenStudy (abhisar):

see the distance or the arc length will remain the same, but the radius will be different now.

OpenStudy (anonymous):

right

OpenStudy (abhisar):

is the radius of the small chain given ?

OpenStudy (abhisar):

is BC given ?

OpenStudy (anonymous):

yes 3 in

OpenStudy (abhisar):

perfect !!

OpenStudy (abhisar):

Now let the angle be x, then x*2\(\pi\)*3=5\(\pi\)/3

OpenStudy (anonymous):

ok let me calculate tht

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

ok i go tht x=13.6isthat correct?

OpenStudy (abhisar):

how ?

OpenStudy (anonymous):

the equation.....

OpenStudy (abhisar):

6\(\pi\)x=5\(\pi\)/3

OpenStudy (anonymous):

6pi x=5.23

OpenStudy (abhisar):

x= \(\huge\frac{5}{18}\)

OpenStudy (abhisar):

getting it ?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

no

OpenStudy (abhisar):

\(\pi\) on Lhs will cancel \(\pi\) on rhs..ok ?

OpenStudy (anonymous):

ok

OpenStudy (abhisar):

okay, other way 6\(\pi\)x=5.2

OpenStudy (anonymous):

yea thats what i got

OpenStudy (abhisar):

18.8*x=5.2

OpenStudy (abhisar):

x=5.2/18.8

OpenStudy (abhisar):

x=0.27

OpenStudy (abhisar):

got it now ?

OpenStudy (anonymous):

5.2/18.8=3.6 and 18.8/5.2=0.27 which one??

OpenStudy (abhisar):

5.2/18.8=0.27

OpenStudy (abhisar):

it's just the way u solve linear equations !

OpenStudy (anonymous):

sorry im tired ya i get it now

OpenStudy (abhisar):

no no it's ok..there is one more step remaining..tjis gave us degrees in radians

OpenStudy (abhisar):

Now we will have to convert it into degrees

OpenStudy (anonymous):

ok so we have .27

OpenStudy (abhisar):

*angle in radians

OpenStudy (abhisar):

yes it is 0.27 radians.....multiply it with \[\frac{180}{\pi}\]

OpenStudy (anonymous):

ok

OpenStudy (abhisar):

it will give u 15.9 = 16°

OpenStudy (anonymous):

right thats what i got.... the answer choices are 112º 100º 120º 106º

OpenStudy (anonymous):

srry i didnt give them b4

OpenStudy (abhisar):

oh my bad !!

OpenStudy (abhisar):

is last one was correct ?

OpenStudy (anonymous):

106?

OpenStudy (abhisar):

No i mean if the answer to the last question u posted was correct ?

OpenStudy (anonymous):

oh the other one ..... i wont know untill i answer this one... theni will know

OpenStudy (abhisar):

ok...give me 5 mins

OpenStudy (anonymous):

sure

OpenStudy (abhisar):

it's 100°

OpenStudy (abhisar):

and answer to previous question will be 32.89

OpenStudy (anonymous):

howd you get it?

OpenStudy (abhisar):

nothying i just recalculated all...there was a calculation mistake !

OpenStudy (abhisar):

*nothing

OpenStudy (abhisar):

The process is same.

OpenStudy (anonymous):

oh thank you so much

OpenStudy (abhisar):

wait for 10 mins and i'll show u the whole process

OpenStudy (abhisar):